OpenStudy (anonymous):

Is the energy of a rotating cylindrical spacecraft conserved when a point-like astronaut climbs up a spoke connecting the walls with the center of the cylinder? If so, when I calculate the fractional change in apparent gravity at the walls when the astronaut reaches the middle I get different answers using conservation of angular momentum and conservation of energy

OpenStudy (anonymous):

Apparent gravity \[g'=r \omega ^{2}\] Consv. of energy: \[\frac{ 1 }{ 2 } (M+m) r ^{2} \omega ^{2}=\frac{ 1 }{ 2 } M r ^{2} \omega ^{2}\] Consv. of L: \[ (M+m) r ^{2} \omega= M r ^{2} \omega\] The first one gives \[\frac{ g-g' }{ g }=\frac{ m }{ M }\] The second one: \[\frac{ g-g' }{ g }=\frac{ m ^{2} }{ M ^{2} }+\frac{ 2m }{ M }\] If i did them right of course